The coefficients 'a' and 'b' of a simple regression line are 12 and 1.3 respectively. What will be the estimated value of the dependent variable if the independent variable is equal to 22?


1
Expert's answer
2021-02-28T17:05:38-0500

In a regression equation

"Y=aX+b"

Where Y is the dependent variable and X is the independent variable.


"Y= 12(22)+1.3=265.3"